# This is the title

### v1

@@
Theorem 1.1.
[2, 10, 17, 11, 1, 7, 16, 14] (Grothendieck Inequality) There is a universal constant $K_{G}$ satisfying the following: For every Hilbert space $\mathcal{H}$ and for all $m,n\in\mathbb{N}$, if a scalar matrix $[a_{j,k}]_{1\leq j\leq m,1\leq k\leq n}$ satisfies
@@
We formulate a problem which is p-adic analogue of Theorem 1.1. We first recall the definition of p-adic Hilbert space.
@@
[…]
@@
(p-adic Grothendieck Inequality Problem) Let $\mathbb{K}$ be a non-Archimedean valued field. Whether there is a universal constant $K_{\mathbb{K}}$ satisfying the following property: for every p-adic Hilbert space $\mathcal{X}$ over $\mathbb{K}$, for all $m,n\in\mathbb{N}$, if $[a_{j,k}]_{1\leq j\leq m,1\leq k\leq n}\in\mathbb{M}_{m\times n}(\mathbb{K})$ satisfies
@@
(p-adic Grothendieck Inequality Problem) Let $\mathbb{K}$ be a non-Archimedean valued field. Whether there is a universal constant $K_{\mathbb{K}}$ satisfying the following property: for every p-adic Hilbert space $\mathcal{X}$ over $\mathbb{K}$, for all $m,n\in\mathbb{N}$, for every matrix $[a_{j,k}]_{1\leq j\leq m,1\leq k\leq n}\in\mathbb{M}_{m\times n}(\mathbb{K})$, for all $u_{1},\dots,u_{m},v_{1},\dots,v_{n}\in\mathcal{X}\setminus\{0\}$, for all $s_{1},\dots,s_{m},t_{1},\dots,t_{n}\in\mathbb{K}\setminus\{0\}$, we have
$\max_{1\leq j\leq m,1\leq k\leq n}\frac{1}{\|u_{j}\|\|v_{k}\|}\left|\sum_{j=1}^{m}\sum_{k=1}^{n}a_{j,k}\langle u_{j},v_{k}\rangle\right|\leq$ $K_{\mathbb{K}}\max_{1\leq j\leq m,1\leq k\leq n}\frac{1}{|s_{j}||t_{k}|}\left|\sum_{j=1}^{m}\sum_{k=1}^{n}a_{j,k}s_{j}t_{k}\right|.$
@@
[15, 12] (Johnson-Lindenstrauss Flattening Lemma) Let $M,N\in\mathbb{N}$ and $\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{M}\in\mathbb{R}^{N}$. For each $10<\varepsilon<$ 1, there exists a Lipschitz map $f:\mathbb{R}^{N}\to\mathbb{R}^{m}$ and a real $r>0$ such that
@@
[4, 12] (Johnson-Lindenstrauss Flattening Lemma: Dasgupta-Gupta form) Let $M,N\in\mathbb{N}$ and $\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{M}\in\mathbb{R}^{N}$. Let $10<\varepsilon<$ 1. Choose any natural number $m$ such that
@@
[12, 8] (Johnson-Lindenstrauss Flattening Lemma: matrix form) There is a universal constant $C>0$ satisfying the following. Let $10<\varepsilon<$ 1, $M,N\in\mathbb{N}$ and $\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{M}\in\mathbb{R}^{N}$. For each natural number
@@
(p-adic Johnson-Lindenstrauss Flattening Problem) Let $\mathbb{K}$ be a non-Archimedean valued field. What is the best function $\phi:(0,1)\times\mathbb{N}\to(0,\infty)$ satisfying the following: There is a universal constant $C>0$ (which may depend upon $\mathbb{K}$) satisfying the following. Let $10<\varepsilon<$ 1, $M,N\in\mathbb{N}$ and $\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{M}\in\mathbb{K}^{N}$. For each natural number

### v2

@@
Theorem $1.$
[2, 10, 17, 11, 1, 7, 16, 14] (Grothendieck Inequality) There is a universal constant $K_{G}$ satisfying the following: For every Hilbert space $\mathcal{H}$ and for all $m,n\in\mathbb{N}$, if a scalar matrix $a_{j,k}]_{1\leq j\leq m,1\leq k\leq n}$ satisfies
@@
We formulate a problem which is p-adic analogue of Theorem $1.$ We first recall the definition of p-adic Hilbert space.
@@
[5, 6, 13] Let $\mathbb{K}$ be a non-Archimedean valued field (with valuation $|\cdot|$) and $\mathcal{X}$ be a non-Archimedean Banach space (with norm $\|\cdot\|$) over $\mathbb{K}$. We say that $\mathcal{X}$ is a p-adic Hilbert space if there is a map (called as p-adic inner product) $\langle\cdot,\cdot\rangle:\mathcal{X}\times\mathcal{X}\to\mathbb{K}$ satisfying following.
@@
(p-adic Grothendieck Inequality Problem) Let $\mathbb{K}$ be a non-Archimedean valued field. Whether there is a universal constant $K_{\mathbb{K}}$ satisfying the following property: for every p-adic Hilbert space $\mathcal{X}$ over $\mathbb{K}$, for all $m,n\in\mathbb{N}$, if $a_{j,k}]_{1\leq j\leq m,1\leq k\leq n}\in\mathbb{M}_{m\times n}(\mathbb{K})$ satisfies
@@
(p-adic Grothendieck Inequality Problem) Let $\mathbb{K}$ be a non-Archimedean valued field. Whether there is a universal constant $K_{\mathbb{K}}$ satisfying the following property: for every p-adic Hilbert space $\mathcal{X}$ over $\mathbb{K}$, for all $m,n\in\mathbb{N}$, for every matrix $a_{j,k}]_{1\leq j\leq m,1\leq k\leq n}\in\mathbb{M}_{m\times n}(\mathbb{K})$, we have
$\sup\left\{\left|\sum_{j=1}^{m}\sum_{k=1}^{n}a_{j,k}\left\langle\frac{u_{j}}{\|u_{j}\|},\frac{v_{k}}{\|v_{k}\|}\right\rangle\right|:u_{1},\dots,u_{m},v_{1},\dots,v_{n}\in\mathcal{X}\setminus\{0\}\right\}\leq$
$K_{\mathbb{K}}\sup\left\{\left|\sum_{j=1}^{m}\sum_{k=1}^{n}a_{j,k}\frac{s_{j}}{|s_{j}|}\frac{t_{k}}{|t_{k}|}\right|:s_{1},\dots,s_{m},t_{1},\dots,t_{n}\in\mathbb{K}\setminus\{0\}\right\}.$
@@
[15, 12] (Johnson-Lindenstrauss Flattening Lemma) Let $M,N\in\mathbb{N}$ and $\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{M}\in\mathbb{R}^{N}$. For each $0<\varepsilon<1$, there exists a Lipschitz map $f:\mathbb{R}^{N}\to\mathbb{R}^{m}$ and a real $r>0$ such that
@@
[4, 12] (Johnson-Lindenstrauss Flattening Lemma: Dasgupta-Gupta form) Let $M,N\in\mathbb{N}$ and $\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{M}\in\mathbb{R}^{N}$. Let $0<\varepsilon<1$. Choose any natural number $m$ such that
@@
[12, 8] (Johnson-Lindenstrauss Flattening Lemma: matrix form) There is a universal constant $C>0$ satisfying the following. Let $0<\varepsilon<1$, $M,N\in\mathbb{N}$ and $\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{M}\in\mathbb{R}^{N}$. For each natural number
@@
(p-adic Johnson-Lindenstrauss Flattening Problem) Let $\mathbb{K}$ be a non-Archimedean valued field. What is the best function $\phi:(0,1)\times\mathbb{N}\to(0,\infty)$ satisfying the following: There is a universal constant $C>0$ (which may depend upon $\mathbb{K}$) satisfying the following. Let $0<\varepsilon<1$, $M,N\in\mathbb{N}$ and $\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{M}\in\mathbb{K}^{N}$. For each natural number